OpenStudy (anonymous):

Its a 7th grade ques, very easy. Please help Construct an equation from the given information and then solve it : I put the same number of chocolate buttons on each of seven cakes. I had 40 buttons at the start and now have 5 buttons left over. How many did i put on each cake ?

OpenStudy (anonymous):

i do know that the answer is 5 but construct an equation and explain.

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

plz help

OpenStudy (baru):

ok, let x be the number of chocolate buttons on each cake

OpenStudy (anonymous):

yes

OpenStudy (baru):

you have 7 cakes, so total number of chocolate buttons=\(7 \times x\) ok?

OpenStudy (anonymous):

yeah

OpenStudy (baru):

i meant total chocolate buttons on the cake

OpenStudy (baru):

but there were still 5 buttons left over, so the total number of buttons 7x+5 (7x on the cake, 5 which are left over)

OpenStudy (baru):

ok?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

then ?

OpenStudy (baru):

the question says u had 40 buttons in total so 7x+5=40

OpenStudy (baru):

ok?

OpenStudy (anonymous):

yeah

OpenStudy (baru):

so this is your equation 7x+5=40

OpenStudy (baru):

can u solve it?

OpenStudy (anonymous):

7x = 40-5 7x=35 x=35/7 x=5

OpenStudy (baru):

yep thats right :)

OpenStudy (baru):

@American.Eagle got it?

OpenStudy (baru):

any question?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

thanks a lot